Question 4

  1. Briefly discuss the advantages and main reasons for leasing.

Consider the following data about a luxury automobile you wish to buy.

  • Current purchase price of the car = $60000
  • Discount rate on loan = 5.0%

Purchase option:

  • Down payment = 20% of purchase price
  • Expected resale value after 3 years = $40000
  • Annual opportunity cost of funds = 8%

  1. If the maturity of the loan is 3 years, what is the monthly payment?
  2. What is the net cost of purchasing the vehicle?

Leasing Option:

The dealer is also offering a 3-year lease with monthly payment of $900, for no money down.
Explain the concept of equity insertion by the lessor?

  1. What is the equity insertion?
  2. What is the lease rate factor?
  3. What is the cost of leasing?
  4. What is the leasing company’s profit and return on the equity investment?
  5. Which option – purchase versus leasing, is better from a strictly financial perspective? Are there any other considerations in the analyses?
Answer #1
a.Advantages and main reasons for leasing.
1.No initial investment
2. No burden of heavy maintenance expenses
3. Conserve your own funds & put to alternative profitable use.
4. Can change the asset more frequently, than owning.
5. Comparaitvely lower monthly rents
6. Can claim tax benefits of lease payments.
b. Monthly payment on the loan
Using the formula to find PV of loan:
PV of loan=Mthly.Pmt.*(1-(1+r)^-n)/r
where   r = 5%/12=0.004167 p.m. & n= 3*12=36 mths.
Loan amt.=60000*(1-20%)= $ 48000
so, 48000=Mthly pmt.*(1-1.004167^-36)/0.004167
Solving the above, we get the monthly pmt. As
48000/((1-1.004167^-36)/0.004167)=
1438.61
c. Net cost of purchasing the vehicle=
Down pmt.+PV of loan=Current purchase price
Current purchase price+Opp/cost of funds locked up-PV of resale price
ie.60000+(60000*8%/12*36)-(40000/1.004167^36)=
39961.36
e. equity insertion is providing own capital, thereby reducing debt .
f. the lease factor here is:
(1-1.004167^-36)/0.004167=
33.37
g. Cost of leasing=
900*33.3655=
30028.95
h.Leasing company's profit=after-tax lease rentals
900*(1-Tax rate)
Return on equity investment=
Profit/Investment in asset
i. Leasing is cheaper
the continuous availability of the automobile
the automobile not becoming out- of -date
are some other concerns that need to be addressed.